# Safe Model-Based Multi-Agent Mean-Field Reinforcement Learning

### v1

@@ C.3.2 Mean-Field Deep Deterministic Policy Gradient (MF-DDPG)
[…]
@@
where $\tilde{\mu}_{n,0}=\mu_{0}$. The MF-MDP formulation in Section C.3.2 turns the Mean-Field Control problem in Section 4 into a Markov Decision Process on the state space of $\mathcal{P}(\mathcal{S})$ and action space $\{\pi:\mathcal{S}\times\mathcal{P}(\mathcal{S})\to\mathcal{A}\}$ with deterministic transition function $U(\cdot)$. We define the Q-value as follows:
@@
Notice that Algorithm 3 uses a model-free approach to optimize the objective, which makes the exploration of particular importance. In comparison to traditional MDPs, MF-MDPs usually have a very constrained set of highly rewarding distributions, and most distributions offer poor rewards, which makes the exploration even more important. This is further complicated by what randomized actions imply in this scenario. In a traditional MDP, we can often assume, for instance, that executing random actions for a fixed number of initial steps would help in finding diversity in the reward space. This is not the case in MF-MDPs – depending on the granularity of the discretization that we use to represent probability distributions, we can expect the mean-field distribution to stay fairly stable. Similarly to [9], we might alleviate this issue by Gaussian mean-field initialization in each episode. Note, however, that this may not necessarily be appropriate in safety-constrained settings, as the initial mean-field distribution is expected to be safe. On top of that, we can add exploration noise to the actions obtained via the policy. Alternatively, we could add noise to the parameters of the policy network for a more consistent approach as in [44].
@@
1:Safety constraint $h_{C}(\cdot)$, calibrated dynamics $\tilde{f}_{n-1}$ represented by $\bm{m}_{n-1}(\cdot)$ and $\bm{\Sigma}_{n-1}(\cdot)$, initial mean-field distribution $\mu_{0}$, known expected reward $\hat{r}(\cdot)$, constants $C_{n,t}$; safety Lipschitz constant $L_{h}$, hyperparameter $\lambda$, number of epochs $K$, number of rollout steps $T$, mini-batch size $B$, learning rate $\alpha$
@@
10:         if $t=0$ then
11:              $c_{t}\leftarrow\hat{r}(\tilde{\mu}_{t},\pi^{\psi}_{t})$
12:         end if
13:         if $11\leq t\leq T-$ 1 then
14:              $c_{t}\leftarrow\hat{r}(\tilde{\mu}_{t},\pi^{\psi}_{t})+\lambda\log(h_{C}(\tilde{\mu}_{t})-L_{h}C_{n,t})$
15:         end if
16:         if $t=T$ then
17:              $c_{t}\leftarrow\lambda\log(h_{C}(\tilde{\mu}_{t})-L_{h}C_{n,t})$
18:         end if
19:         $R\leftarrow R\cup\{(\tilde{\mu}_{t},c_{t},\tilde{\mu}_{t+1})\}$
20:         Sample a mini-batch of $B$ random transitions $\{(\tilde{\mu}_{i},c_{i},\tilde{\mu}_{i+1})\}_{i=1}^{B}$ from $R$
21:         for $i=1,\ldots,B$ do
22:              $q_{i}\leftarrow c_{i}+\gamma Q^{\theta^{\prime}}_{t}(\tilde{\mu}_{i+1},\bm{\pi}^{\psi^{\prime}}(\tilde{\mu}_{i+1}))$
23:         end for
24:         Update $\theta$ with gradient descent $\nabla_{\theta}\frac{1}{B}\sum_{i}(q_{i}-Q^{\theta}_{t}(\tilde{\mu}_{i},\bm{\pi}^{\psi}(\tilde{\mu}_{i}))^{2}$
25:         Update $\psi$ with gradient ascent $\nabla_{\psi}\frac{1}{B}\sum_{i}Q^{\theta}_{t}(\tilde{\mu}_{i},\bm{\pi}^{\psi}(\tilde{\mu}_{i}))$
26:         $\theta^{\prime}\leftarrow\alpha\theta+(1-\alpha)\theta^{\prime}$
27:         $\psi^{\prime}\leftarrow\alpha\psi+(1-\alpha)\psi^{\prime}$
28:     end for
29:end for
30:$\bm{\pi}_{n}^{\psi}\leftarrow\bm{\pi}^{\psi^{\prime}}$
Appendix D Experiments
In this section, we provide further analysis, the motivation behind our modeling decisions, and details for making our experiments easily replicable. We use a private cluster with GPUs to run our experiments. Safe-$\text{M}^{3}$-UCRL and $\text{M}^{3}$-UCRL with known dynamics each used 15 minutes of one Intel Xeon Gold 511 CPU core, 32 GB of RAM and one Nvidia GeForce RTX 3090 GPU. Training Safe-$\text{M}^{3}$-UCRL and $\text{M}^{3}$-UCRL with unknown dynamics to produce results in Figure 3 and Figure 4(b) had 24 hours of access to fifty Intel Xeon Gold 511 CPU cores, 64 GB of RAM and fifty Nvidia GeForce RTX 3090 GPUs during the batch execution necessary for training. The evaluation, i.e., generating the results for, e.g., Figure 5(b) and Figure 7 took around 1 hour of one Intel Xeon Gold 511 CPU core, 64 GB of RAM and one Nvidia GeForce RTX 3090 GPU. The only computationally intensive evaluation task was for Figure 5(a) for more than 1 million agents. We had access to Xeon Gold 511 CPU core, 128-256 GB of RAM, for around 8 hours. The implementation was predominantly done in Python packages PyTorch [42] and NumPy [24].
@@ Appendix D Experiments
    • Estimating the demand distribution for taxi service $\rho_{0}$
@@
    • Learning unknown transition dynamics using probabilistic ensemble, i.e., statistical estimators $\bm{m}_{n-1}(\cdot)$ and $\bm{\Sigma}_{n-1}(\cdot)$
  4. 4.
@@
D.1 Input Data Preprocessing
@@
We consider taxi trajectories collected in Shenzhen’s extended city center with the geographical area spanning from 114.015 to 114.14 degrees longitude and from 22.5 to 22.625 degrees latitude. We have access to the trajectories of five full working weeks (Monday to Sunday) collected between 18th January 2016 and 25th September 2016. We restrict ourselves to evening peak hours between 19:00 and 22:00. We represent probability distributions by discretizing the space into $k\times k$ unit grid with $k=25$ where each cell, $C_{ij}=[\frac{i}{k},\frac{i+1}{k}\rangle\times[\frac{j}{k},\frac{j+1}{k}\rangle$ for $i,j\in\{0,\ldots,k-1\}$, represents a square neighborhood of around $550\times 550$ meters on the city map. We represent taxi service demand distribution $\rho_{0}\in\mathcal{P}(\mathcal{S})$ as a $k\times k$ matrix where entries $[\rho_{0}]_{ij},i,j\in\{0,\ldots,k-1\}$ represent a probability of a trip originating in the neighborhood $C_{ij}$. The probabilities are estimated as an average over the considered period and kept constant during each step $t$ of the learning protocol (see Section D.3). To smooth out the demand distribution and remove possible noise in the raw data, we apply 2-dimensional median smoothing with a window equal to 3 and show the output in Figure 1(a). If the passenger’s trip originates at the state $s\in\mathcal{S}$ the likelihood of its destinations is defined by the mapping $\Phi:\mathcal{S}\rightarrow\mathcal{P}(\mathcal{S})$ which we use in Section D.2 to define sequential transitions by first executing passenger trips followed by vehicle repositioning (see Figure 1). We flatten the $k\times k$ space grid into $k^{2}$-dimensional vector and represent $\Phi(\cdot)$ as a $k^{2}\times k^{2}$ matrix. The entries $[\Phi]_{ij}$ with $i,j\in\{0,\ldots,k^{2}-1\}$ represent the likelihood of a passenger’s trip that originated in the neighborhood $i$ ending in the neighborhood $j$. The likelihoods are estimated as an average over the considered period and kept constant. Figure 4 shows $\Phi$ column average, i.e., the trip destinations’ likelihood given an arbitrary trip origin.
@@
Figure 4: Passenger’s trip destinations’ likelihood during evening peak hours. The image shows the trip destinations’ likelihood given an arbitrary trip origin (see Section D.1). We observe that the trips to most residential areas are almost equally likely and dispersed across the entire study region, which is consistent with the intuition that some residents commute back home (outside of the city center) after a work day, while others commute towards the city center for, e.g., leisure activities.
D.2 Modeling Assumptions and Model Parameters
@@
We represent our area of interest as a two-dimensional unit square, i.e., the state-space $\mathcal{S}=[0,1]^{2}$, and assume that taxis can move freely using repositioning actions taken from $\mathcal{A}=[-1,1]^{2}$. If the action takes a taxi outside of the state-space borders, we project it back onto the border. Since the fleet is operating in a noisy traffic environment, repositioning usually cannot be executed perfectly. We model the noise $\varepsilon_{n,t}\sim\text{TN}(0,\sigma^{2}I_{2})$ by a Gaussian with a known variance $\sigma^{2}$ truncated at the borders of $\mathcal{S}$ and $I_{2}$ is the $2\times 2$ unit matrix. We use a standard deviation $\sigma=0.0175$ to represent that the passenger will be dropped off with a 68% probability within a circle with a 240-meter radius of the requested destination or with 95% probability within a circle with a 480-meter radius. For simplicity, we assume that every passenger ride lasts fifteen minutes and that the repositioning is executed instantaneously. We assume that the representative agent (or multiple representative agents) reports the trajectories obtained during the interaction with the environment to the global controller at the end of the day. Therefore, we discretize our modeling horizon in fifteen-minute operational intervals, i.e., $T=12$, and each episode $n$ represents one day. The goal of the service provider is maximizing the profit, which correlates with the amount of satisfied demand. Therefore, we implicitly model the profit by the negative of the Kullback-Leibler divergence between the taxis’ distribution $\mu_{n,t}$ and demand for taxi service $\rho_{0}$, i.e., $r(s_{n,t},\mu_{n,t},a_{n,t})=-D_{KL}(\rho_{0}||\mu_{n,t})$. In other words, $r(\cdot)$ measures the similarity between taxis’ and demand distributions; the closer the distributions higher the profit. In practice, greedy profit maximization is often prevented by imposing service accessibility requirements by regulatory bodies. We use entropic safety constraint $h_{C}(\mu_{n,t})=-\int_{\mathcal{S}}\log(\mu_{n,t}(s))\mu_{n,t}(ds)-C$ to enforce the service accessibility across all residential areas. Therefore, the optimization objective in Section 4 trades off between maximizing the profit $r(\cdot)$ and adhering to accessibility requirements imposed by $h_{C}(\cdot)$. We use matrix $\Phi$ introduced Section D.1 to model sequential transitions by first executing passenger trips followed by vehicle repositioning. Formally, the unknown system dynamics are given by $f(s_{n,t},\mu_{n,t},a_{n,t})=\text{clip}(s^{\Phi}_{n,t}+a_{n,t},0,1)+\varepsilon_{n,t}$, where $s^{\Phi}_{n,t}=\text{sample}(\Phi(s_{n,t}))$. The sampling is done by first determining the destination cell $C_{ij}$ given the origin cell determined by the current state $s_{n,t}$ and origin-destination likelihood matrix $\Phi(\cdot)$. Secondly, we determine the destination state $s^{\Phi}_{n,t}$ by uniform sampling from a cell $C_{ij}$, which is a simplified model of the passenger preferences of the final destination.
@@
Mean-field transition dynamics. During the learning/training phase, we assume that the number of agents $m\rightarrow\infty$ and that these agents induce the mean-field distribution $\mu_{n,t}$ for episode $n$ and step $t$. But, one of the major practical challenges is implementing the mean-field transition function $U(\cdot)$ (see Equation 2). The main difficulties are representing the mean-field distribution $\mu_{n,t}$ and computing the integral in Equation 2 over the state space $\mathcal{S}$. We use discretization to represent the mean-field distribution even though other representations, such as a mixture of Gaussians, are possible. Concretely, we represent the mean-field distribution as $\mu_{n,t}=[\mu_{n,t}]_{ij}$ with $i,j\in\{0,\ldots,k-1\}$ by associating the probability $[\mu_{n,t}]_{ij}=\mathbb{P}\big{[}s_{n,t}\in C_{ij}\big{]}$ of the representative agent residing within each cell $C_{ij}$ during episode $n$ at step $t$. The initial mean-field distribution, $\mu_{n,0}$, follows the uniform distribution to maximize the Shannon entropy, ensuring the safety at the beginning of every episode $n$ at $t=0$, i.e., $h_{C}(\mu_{n,0})\geq 0$. In vehicle repositioning, the mean-field transition dynamics $U(\cdot)$ consist of two sequential steps induced by steps in individual system dynamics $f$. Namely, first, the demand shifts the mean-field distribution, followed by the transition induced by the controller. Formally, the mean-field demand transition is computed as $\mu_{n,t}^{\Phi}=(\mu_{n,t}\cdot p)\times\Phi+\mu_{n,t}\cdot(1-p)$, where $\cdot$ denotes elementwise multiplication, $\times$ denotes matrix multiplication and $p=\min(1,\frac{\rho_{0}}{\mu_{n,t}})$ represents a proportion of occupied taxis. The mean-field controller transition requires computing the integral in Equation 2 for which we use the discrete approximation given the points $c_{ij}$ uniformly chosen from cells $C_{ij}$ for $i,j\in\{0,\ldots,k-1\}$:
$[\mu_{n,t+1}]_{ij}=\sum_{k,l}\mathbb{P}[f(c_{kl},\mu_{n,t}^{\Phi},\pi_{n,t}(c_{kl},\mu_{n,t}^{\Phi}))+\varepsilon_{n,t}\in C_{ij}][\mu_{n,t}^{\Phi}]_{kl},$
@@
for the episode $n$ and step $t$. We assume that the noise term $\varepsilon_{n,t}$ is independent across episodes and steps as well as along the two dimensions while the truncation parameters are adjusted relative to the state space borders. Thus, we have the following:
@@
$\begin{split}&\mathbb{P}\left[f(c_{kl},\mu^{\Phi}_{n,t},\pi_{n,t}(c_{kl},\mu_{n,t}^{\Phi}))+\varepsilon_{n,t}\in C_{ij}\right]\\ &=\mathbb{P}\left[f(c_{kl},\mu^{\Phi}_{n,t},\pi_{n,t}(c_{kl},\mu_{n,t}^{\Phi}))_{x}+\varepsilon_{n,t,x}\in\left[\frac{i}{k},\frac{i+1}{k}\right\rangle\right]\\ &\quad\times\mathbb{P}\left[f(c_{kl},\mu^{\Phi}_{n,t},\pi_{n,t}(c_{kl},\mu_{n,t}^{\Phi}))_{y}+\varepsilon_{n,t,y}\in\left[\frac{j}{k},\frac{j+1}{k}\right\rangle\right]\\ &=\left[\phi\left(\frac{i+1}{k}-f(c_{kl},\mu^{\Phi}_{n,t},\pi_{n,t}(c_{kl},\mu_{n,t}^{\Phi}))_{x}\right)-\phi\left(\frac{i}{k}-f(c_{kl},\mu^{\Phi}_{n,t},\pi_{n,t}(c_{kl},\mu_{n,t}^{\Phi}))_{x}\right)\right]\\ &\quad\cdot\left[\phi\left(\frac{j+1}{k}-f(c_{kl},\mu^{\Phi}_{n,t},\pi_{n,t}(c_{kl},\mu_{n,t}^{\Phi}))_{y}\right)-\phi\left(\frac{j}{k}-f(c_{kl},\mu^{\Phi}_{n,t},\pi_{n,t}(c_{kl},\mu_{n,t}^{\Phi}))_{y}\right)\right],\end{split}$
@@
Mean-field transition dynamics in the finite regime In Section D.5, we instantiate a finite number of taxis $m<\infty$ to evaluate the policy performance in a realistic setting. We keep track of taxis’ states $s_{t}^{(l)}$ for every taxi $l\in\{1,\ldots,m\}$ at steps $t=0,\ldots,T$. In this setting, we approximate the mean-field transition $U(\cdot)$ with the normalized two-dimensional histogram $[\mu_{t+1}]_{ij}$ with bins defined by the cells $C_{ij}$ for $i,j\in\{0,\ldots,k-1\}$ given taxis’ next states $s_{t+1}^{(l)}=f(s_{t}^{(l)},\mu_{t}^{\Phi},\pi(s_{t}^{(l)},\mu_{t}^{\Phi}))+\varepsilon_{t}$.
D.3 Model-based Learning Protocol

### v2

@@ C.3.2. Mean-Field Deep Deterministic Policy Gradient (MF-DDPG)
subject $\text{subject to}\quad\tilde{f}_{n-1}(\tilde{z})$ ${}={}\bm{m}_{n-1}(\tilde{z}){}+{}\beta_{n-1}\bm{\Sigma}_{n-1}(\tilde{z})\eta^{\psi}(\tilde{z})$ (12b)
@@
where $\tilde{\mu}_{n,0}{}={}\mu_{0}$ for every $n$. The MF-MDP formulation in Section C.3.2 turns the MFC in Section 4 into a Markov Decision Process on the state space of $\mathcal{P}(\mathcal{S})$ and action space $\{\pi:\mathcal{S}\times\mathcal{P}(\mathcal{S})\to\mathcal{A}\}$ with deterministic transition function $U(\cdot)$. We define the Q-value as follows
@@
Notice that Algorithm 3 uses a model-free approach to optimize the objective, which makes the exploration of particular importance. In comparison to traditional MDPs, MF-MDPs usually have a very constrained set of highly rewarding distributions, and most distributions offer poor rewards, which makes the exploration even more important. This is further complicated by what randomized actions imply in this scenario. In a traditional MDP, we can often assume, for instance, that executing random actions for a fixed number of initial steps would help in finding diversity in the reward space. This is not the case in MF-MDPs – depending on the granularity of the discretization that we use to represent probability distributions, we can expect the mean-field distribution to stay fairly stable. Similarly to Carmona et al. (2019b), we might alleviate this issue by Gaussian mean-field initialization in each episode. Note, however, that this may not necessarily be appropriate in safety-constrained settings, as the initial mean-field distribution is expected to be safe. On top of that, we can add exploration noise to the actions obtained via the policy. Alternatively, we could add noise to the parameters of the policy network for a more consistent approach as in Plappert et al. (2017).
@@
1:Safety constraint $h_{C}(\cdot)$, calibrated transitions $\tilde{f}_{n-1}$ represented by $\bm{m}_{n-1}(\cdot)$ and $\bm{\Sigma}_{n-1}(\cdot)$, initial mean-field distribution $\mu_{0}$, known expected reward $\hat{r}(\cdot)$, constants $C_{n,t}$; safety Lipschitz constant $L_{h}$, hyperparameter $\lambda$, number of epochs $K$, number of rollout steps $T$, mini-batch size $B$, learning rate $\alpha$
@@
10:         $c_{t}\leftarrow\hat{r}(\tilde{\mu}_{t},\pi^{\psi}_{t}){}+{}\lambda\log(h_{C}(\tilde{\mu}_{t+1})-L_{h}C_{n,t+1})$
11:         $R\leftarrow R\cup\{(\tilde{\mu}_{t},c_{t},\tilde{\mu}_{t{}+{}1})\}$
12:         Sample a mini-batch of $B$ random transitions $\{(\tilde{\mu}_{i},c_{i},\tilde{\mu}_{i{}+{}1})\}_{i=1}^{B}$ from $R$
13:         for $i=1,\ldots,B$ do
14:              $q_{i}\leftarrow c_{i}{}+{}\gamma Q^{\theta^{\prime}}_{t}(\tilde{\mu}_{i{}+{}1},\bm{\pi}^{\psi^{\prime}}(\cdot,\tilde{\mu}_{i{}+{}1}))$
15:         end for
16:         Update $\theta$ with gradient descent $\nabla_{\theta}\frac{1}{B}\sum_{i}(q_{i}-Q^{\theta}_{t}(\tilde{\mu}_{i},\bm{\pi}^{\psi}(\cdot,\tilde{\mu}_{i}))^{2}$
17:         Update $\psi$ with gradient ascent $\nabla_{\psi}\frac{1}{B}\sum_{i}Q^{\theta}_{t}(\tilde{\mu}_{i},\bm{\pi}^{\psi}(\cdot,\tilde{\mu}_{i}))$
18:         $\theta^{\prime}\leftarrow\alpha\theta{}+{}(1-\alpha)\theta^{\prime}$
19:         $\psi^{\prime}\leftarrow\alpha\psi{}+{}(1-\alpha)\psi^{\prime}$
20:     end for
21:end for
22:$\bm{\pi}_{n}^{\psi}\leftarrow\bm{\pi}^{\psi^{\prime}}$
Appendix D Experiments – Vehicle Repositioning
In this section, we provide further analysis, the motivation behind our modeling decisions, and details for making our experiments easily replicable. We use a private cluster with GPUs to run our experiments. Safe-M3-UCRL and M3-UCRL under known transitions each used 15 minutes of one Intel Xeon Gold 511 CPU core, 32 GB of RAM and one Nvidia GeForce RTX 3090 GPU. Training Safe-M3-UCRL and M3-UCRL under unknown transitions to produce results in Figure 1(d), Figure 1(f) and Figure 7(b) had 24 hours of access to fifty Intel Xeon Gold 511 CPU cores, 64 GB of RAM and fifty Nvidia GeForce RTX 3090 GPUs during the batch execution necessary for training. The evaluation, i.e., generating the results for, e.g., Figure 8(b) and Figure 10 took around 1 hour of one Intel Xeon Gold 511 CPU core, 64 GB of RAM and one Nvidia GeForce RTX 3090 GPU. The only computationally intensive evaluation task was for Figure 8(a) for more than 1 million agents. We had access to Xeon Gold 511 CPU core, 128-256 GB of RAM, for around 8 hours. The implementation was predominantly done in Python packages PyTorch Paszke et al. (2017) and NumPy Harris et al. (2020).
@@ Appendix D Experiments – Vehicle Repositioning
    • Estimating the demand distribution for the service $\rho_{0}$
@@
    • Learning unknown transitions using probabilistic neural network ensemble, i.e., statistical estimators $\bm{m}_{n-1}(\cdot)$ and $\bm{\Sigma}_{n-1}(\cdot)$
  4. (4)
@@
D.1. Input Data Preprocessing
@@
We consider vehicle trajectories collected in Shenzhen’s extended city center with the geographical area spanning from 114.015 to 114.14 degrees longitude and from 22.5 to 22.625 degrees latitude. We have access to the trajectories of five full working weeks (Monday to Sunday) collected between 18th January 2016 and 25th September 2016. We restrict ourselves to evening peak hours between 19:00 and 22:00. We represent probability distributions by discretizing the space into $k\times k$ unit grid with $k=25$ where each cell, $C_{ij}=[\frac{i}{k},\frac{i{}+{}1}{k}\rangle\times[\frac{j}{k},\frac{j{}+{}1}{k}\rangle$ for $i,j\in\{0,\ldots,k-1\}$, represents a square neighborhood of around $550\times 550$ meters on the city map. We represent the service demand distribution $\rho_{0}\in\mathcal{P}(\mathcal{S})$ as a $k\times k$ matrix where entries $[\rho_{0}]_{ij},i,j\in\{0,\ldots,k-1\}$ represent a probability of a trip originating in the neighborhood $C_{ij}$. The probabilities are estimated as an average over the considered period and kept constant during each step $t$ of the learning protocol (see Section D.3). To smooth out the demand distribution and remove possible noise in the raw data, we apply 2-dimensional median smoothing with a window equal to 3 and show the output in Figure 2(a). If the passenger’s trip originates at the state $s\in\mathcal{S}$ the likelihood of its destinations is defined by the mapping $\Phi:\mathcal{S}\rightarrow\mathcal{P}(\mathcal{S})$ which we use in Section D.2 to define sequential transitions by first executing passenger trips followed by vehicle repositioning (see Figure 1). We flatten the $k\times k$ space grid into $k^{2}$-dimensional vector and represent $\Phi(\cdot)$ as a $k^{2}\times k^{2}$ probability matrix, i.e., rows summing to one represent outgoing mass for each cell. The entries $[\Phi]_{ij}$ with $i,j\in\{0,\ldots,k^{2}-1\}$ represent the likelihood of a passenger’s trip that originated in the neighborhood $i$ ending in the neighborhood $j$. The likelihoods are estimated as an average over the considered period and kept constant. Figure 4 shows $\Phi$ column average, i.e., the trip destinations’ likelihood given an arbitrary trip origin.
@@
Figure 4. Passenger’s trip destinations’ likelihood during evening peak hours given an arbitrary trip origin (see Section D.1). We observe that the trips to most residential areas are almost equally likely and dispersed across the entire study region, which is consistent with the intuition that some residents commute back home (outside of the city center) after a work day, while others commute towards the city center for, e.g., leisure activities.
Passenger’s trip destinations’ likelihood during evening peak hours given an arbitrary trip origin. We observe that the trips to most residential areas are almost equally likely and dispersed across the entire study region, which is consistent with the intuition that some residents commute back home (outside of the city center) after a work day, while others commute towards the city center for, e.g., leisure activities.
Figure 4. Passenger’s trip destinations’ likelihood during evening peak hours given an arbitrary trip origin (see Section D.1). We observe that the trips to most residential areas are almost equally likely and dispersed across the entire study region, which is consistent with the intuition that some residents commute back home (outside of the city center) after a work day, while others commute towards the city center for, e.g., leisure activities.
D.2. Modeling Assumptions and Model Parameters
@@
We represent our area of interest as a two-dimensional unit square, i.e., the state-space $\mathcal{S}{}={}[0,1]^{2}$, and assume that vehicles can move freely using repositioning actions taken from $\mathcal{A}{}={}[-1,1]^{2}$. If the action takes a vehicle outside of the state-space borders, we project it back onto the border. Since the fleet is operating in a noisy traffic environment, repositioning usually cannot be executed perfectly. We model the noise $\varepsilon_{n,t}\sim\text{TN}(0,\sigma^{2}I_{2})$ by a Gaussian with a known variance $\sigma^{2}$ truncated at the borders of $\mathcal{S}$ and $I_{2}$ is the $2\times 2$ unit matrix. We use a standard deviation $\sigma=0.0175$ to represent that the vehicle will be repositioned with a 68% probability within a circle with a 240-meter radius of the desired location or with 95% probability within a circle with a 480-meter radius. For simplicity, we assume that every passenger ride lasts fifteen minutes and that the repositioning is executed instantaneously. We assume that the representative agent (or multiple representative agents) reports the trajectories obtained during the interaction with the environment to the global controller at the end of the day. Therefore, we discretize our modeling horizon in fifteen-minute operational intervals, i.e., $T=12$, and each episode $n$ represents one day. The goal of the service provider is maximizing the profit, which correlates with the amount of satisfied demand. Therefore, we model the service provider goal of maximizing the coverage of the demand by the negative of the Kullback-Leibler divergence between the vehicles’ distribution $\mu_{n,t}$ and demand for service $\rho_{0}$, i.e., $r(z_{n,t}){}={}-D_{KL}(\rho_{0}||\mu_{n,t})$. In other words, $r(\cdot)$ measures the similarity between vehicles’ and demand distributions; the closer the distributions, the higher the profit. In practice, greedy profit maximization is often prevented by imposing service accessibility requirements by regulatory bodies. Due to the discrete representation of the mean-field distribution (discussed in the next paragraph), we use Shannon entropy to define the safety constraint
$h_{C}(\mu_{n,t}){}={}-\sum_{i,j}\log([\mu_{n,t}]_{ij})[\mu_{n,t}]_{ij}-C$ (13)
to enforce the service accessibility across all residential areas. Therefore, the optimization objective in Section 4 trades off between maximizing the profit $r(\cdot)$ and adhering to accessibility requirements imposed by $h_{C}(\cdot)$. We use matrix $\Phi$ introduced Section D.1 to model sequential transitions by first executing passenger trips followed by vehicle repositioning. Formally, the next state $s_{n,t{}+{}1}{}={}f(z_{n,t}){}+{}\varepsilon_{n,t}$ is induced by the unknown transitions $f(z_{n,t}){}={}\text{clip}(s^{\Phi}_{n,t}{}+{}a_{n,t},0,1)$, where $s^{\Phi}_{n,t}\sim\Phi(s_{n,t})$. Firstly, we find origin cell $i\in k^{2}$ such that $s_{n,t}$ resides in it and sample destination cell $j\in k^{2}$ given likelihoods defined in row $i$ of the probability matrix $\Phi$. Secondly, we determine the destination state $s^{\Phi}_{n,t}$ by uniform sampling from the destination cell $j$, which is a simplified model of the passenger preferences of the final destination. Notice that the controller determines repositioning actions given intermediate states $s^{\Phi}_{n,t}$ obtained after executing passenger trips.
@@
Mean-field transitions. During the learning/training phase, we assume that the number of agents $m\rightarrow\infty$ and that these agents induce the mean-field distribution $\mu_{n,t}$ for episode $n$ and step $t$. But, one of the major practical challenges is implementing the mean-field transition function $U(\cdot)$ (see Equation 2). The main difficulties are representing the mean-field distribution $\mu_{n,t}$ and computing the integral in Equation 2. We use discretization to represent the mean-field distribution even though other representations, such as a mixture of Gaussians, are possible. Concretely, we represent the mean-field distribution as $\mu_{n,t}{}={}[\mu_{n,t}]_{ij}$ with $i,j\in\{0,\ldots,k-1\}$ with $k=25$ by associating the probability $[\mu_{n,t}]_{ij}{}={}\mathbb{P}\big{[}s_{n,t}\in C_{ij}\big{]}$ of the representative agent residing within each cell $C_{ij}$ during episode $n$ at step $t$. Note that the discrete representation of the mean-field distribution does not affect the state and action spaces, which remain continuous. The initial mean-field distributions, $\mu_{n,0}$ for every $n$, follow the uniform distribution which maximizes the Shannon entropy, ensuring the safety at the beginning of every episode $n$ at $t=0$, i.e., $h_{C}(\mu_{n,0}){}\geq{}0$. In vehicle repositioning, the mean-field transitions $U(\cdot)$ consist of two sequential steps induced by transitions $f$. Namely, first, the demand shifts the mean-field distribution, followed by the transition induced by the controller. Formally, the mean-field demand transition is computed as $\mu_{n,t}^{\Phi}=(\mu_{n,t}\cdot p)\times\Phi{}+{}\mu_{n,t}\cdot(1-p)$, where $\cdot$ denotes elementwise multiplication, $\times$ denotes matrix multiplication and $p=\min(1,\frac{\rho_{0}}{\mu_{n,t}})$ represents elementwise proportion of occupied vehicles. The mean-field controller transition requires computing the integral in Equation 2 for which we use the discrete approximation given the points $c_{ij}$ uniformly chosen from cells $C_{ij}$ for $i,j\in\{0,\ldots,k-1\}$
$[\mu_{n,t{}+{}1}]_{ij}{}={}\sum_{k,l}\mathbb{P}[f(c_{kl},\mu_{n,t}^{\Phi},\pi_{n,t}(c_{kl},\mu_{n,t}^{\Phi})){}+{}\varepsilon_{n,t}\in C_{ij}][\mu_{n,t}^{\Phi}]_{kl},$ (14)
@@
for the episode $n$ and step $t$. We assume that the noise term $\varepsilon_{n,t}$ is independent across episodes and steps as well as along the two dimensions while the truncation parameters are adjusted relative to the state space borders. Thus, we have the following
@@
$\begin{split}&\mathbb{P}\left[f(c_{kl},\mu^{\Phi}_{n,t},\pi_{n,t}(c_{kl},\mu_{n,t}^{\Phi})){}+{}\varepsilon_{n,t}\in C_{ij}\right]\\ ={}&\mathbb{P}\left[f(c_{kl},\mu^{\Phi}_{n,t},\pi_{n,t}(c_{kl},\mu_{n,t}^{\Phi}))_{x}{}+{}\varepsilon_{n,t,x}\in\left[\frac{i}{k},\frac{i{}+{}1}{k}\right\rangle\right]\\ &\times\mathbb{P}\left[f(c_{kl},\mu^{\Phi}_{n,t},\pi_{n,t}(c_{kl},\mu_{n,t}^{\Phi}))_{y}{}+{}\varepsilon_{n,t,y}\in\left[\frac{j}{k},\frac{j{}+{}1}{k}\right\rangle\right]\\ ={}&\left[\phi\left(\frac{i{}+{}1}{k}-f(c_{kl},\mu^{\Phi}_{n,t},\pi_{n,t}(c_{kl},\mu_{n,t}^{\Phi}))_{x}\right)\right.\\ &\left.-\;\phi\left(\frac{i}{k}-f(c_{kl},\mu^{\Phi}_{n,t},\pi_{n,t}(c_{kl},\mu_{n,t}^{\Phi}))_{x}\right)\right]\\ &\cdot\left[\phi\left(\frac{j{}+{}1}{k}-f(c_{kl},\mu^{\Phi}_{n,t},\pi_{n,t}(c_{kl},\mu_{n,t}^{\Phi}))_{y}\right)\right.\\ &\left.-\;\phi\left(\frac{j}{k}-f(c_{kl},\mu^{\Phi}_{n,t},\pi_{n,t}(c_{kl},\mu_{n,t}^{\Phi}))_{y}\right)\right],\end{split}$ (15)
@@
Mean-field transitions in the finite regime In Section D.5, we instantiate a finite number of vehicles $m<\infty$ to evaluate the policy performance in a realistic setting. We keep track of vehicles’ states $s_{t}^{(l)}$ for every vehicle $l\in\{1,\ldots,m\}$ at steps $t{}={}0,\ldots,T$. In this setting, we approximate the mean-field transition $U(\cdot)$ with the normalized two-dimensional histogram $[\mu_{t{}+{}1}]_{ij}$ with bins defined by the cells $C_{ij}$ for $i,j\in\{0,\ldots,k-1\}$ given vehicles’ next states $s_{t{}+{}1}^{(l)}{}={}f(s_{t}^{(l)},\mu_{t}^{\Phi},\pi(s_{t}^{(l)},\mu_{t}^{\Phi})){}+{}\varepsilon_{t}$.
D.3. Model-Based Learning Protocol